### v1

@@ 2.4 Case study: GPCR metabolite and drug target binding and identification
We fine-tune our multi-view model on experimental binding affinity assays for 106 GPCR targets. This dataset contains experimental pKI values curated from ChEMBL [36] and GLASS [37]. GPCRs with fewer bioactivity measurements ($<100$) were not considered. Before training, the dataset is filtered to exclude molecules with a molecular weight $>600$ Da. This cutoff is commiserate with the molecular properties of the metabolites and drug molecules that are in the screening (inference) set. Datasets with less than 50 data points were removed from the set. Each assay/GPCR target is treated as a separate regression task and data is split using the balanced scaffold protocol [25]. In Fig 3A, we plot the prediction on the held-out set against the experimental values. The multi-view fine-tuned models are overall in good agreement with the held-out set (Pearson correlation coefficient 0.78, average RMSE 0.79). To further validate our model, in Fig. 3B we plot the Pearson correlation coefficient against the RMSE for each individual target. These metrics are negatively correlated, as expected. Over $80\%$ of fine-tuned models have $\text{RMSE}<1$ and nearly $70\%$ have correlations $>0.6$ on the held-out set.
@@
Next we consider interactions of molecules in our virtual screen with specific AD-related GPCRs. FPR1 is suggested as a genetics-informed GPCR. High level of FPR1 is causally associated with increased risks of AD (Beta = -0.1, FDR = 0.001, Fig. S5A). Through inspecting top 10 small molecules in our screen, we found that a gut metabolite, acetyl-glutamine is predicted to interact with FPR1 (multi-view score: 6.8). By re-analyzing the abundance of gut metabolites in microbial strains in vitro [51], acetyl-glutamine was determined to have a high level in Ruminococcus gnavus (log2FoldChange: 1.19), a specie that mono-colonized in mice performed better on a spatial working memory test [52]. Additionally, over-expression of $\text{A}\beta$ protein is significantly characterized by the decreased Ruminococcus gnavus [53]. Three-dimension (3D) binding interaction analysis revealed that acetyl-glutamine is located in an allosteric binding site that is distinct from the classical binding site, indicating a potential regulation mechanism (Fig. 3C). For top predicted drugs, glutathione (GSH), an anti-oxidant drug [54], GSH, is prioritized (multi-view score: 6.53) to interact with FPR1. GSH is a nutrition supplementation that consist of three amino acids, including glutamic acid [55]. The depletion of GSH has been found in the hippocampal regions in mild cognitive impairment (MCI) and AD patients compared to healthy control [56]. GSH forms strong hydrogen bond interactions with FPR1 in the classic binding site (Fig. 3D). We also generated the attention heatmap of small molecules. We found that most of functional groups of acetyl-glutamine (e.g., amide group) and GSH have higher attentions (Fig. 3C-D).
@@
ADA2A is a GPCR with the strong multi-omics evidence to have a significant lower expression in AD (differential expressed in 1 human and 4 mouse transcriptomics datasets, Fig. S5B) [50]. A gut metabolite, fructose 1,6-biphosphate, is predicted to strongly interact with ADA2A (multi-view score 7.37). Fructose 1,6-biphosphate is involved in glycolysis, the level of which was down regulated in APP/PS1 mice [57]. High levels of 1,6-biphosphate have been detected in genus Mitsuokella (log2FoldChange: 13.42), which is decreased in $\text{A}\beta+$ (positive) patients[58]. Fructose 1,6-biphosphate is predicted to bind a classic site of ADA2A (Fig. 3E). Isosorbide dinitrate, a drug used to prevent and treat chest pain caused by coronary artery disease[38], is predicted as a top candidate to bind ADA2A (multi-view score: 8.21, Fig. 3F). Similarly, the high attention of fructose 1,6-biphosphate and Isosorbide dinitrate were captured for the key chemical moieties (Fig. 3E-F).
@@
Figure 3: (A) Correlation plot of predicted vs. experimental pKI values in the held out set. Values of the Pearson correlation and the RMSE are given in the legend. (B) Pearson’s correlation coefficient plotted against RMSE for 33 Alzheimer’s disease related GPCR targets (red dots) and others (blue dots). Examples of strong binders to FPR1 (C-D) and ADA2A (E-F) are depicted in the respective panels. Binding poses generated from docking software are depicted alongside heat maps on attention levels as predicted by the CNN sub-network of our model.
@@ 3 Discussion
A molecular foundation model that is fine-tunable on a broad, diverse set of tasks can serve as a crucial component of any discovery workflow. The choice of molecular representation is key to building such foundation models. However, each choice may have strengths and weaknesses, and each individual representation (view) typically performs well on certain downstream tasks and not others. Here we introduce MMELON, a multi-view architecture that combines molecular representations from multiple embeddings using an interpretable late fusion approach, utilizing foundation models based on three common representations of chemical structure, Graph, Image and Text.
We fine-tune our multi-view model on a set benchmark datasets that span diverse tasks such as binding activity, toxicity and drug metabolism. We compare the performance of the multi-view model with the single view models whose encodings feed into it. Overall the multi-view model performs robustly across a wide range of property prediction tasks involving both classification and regression - it gives the best performance in the majority of the tasks, while closely matching the best single model in the remaining ones. In contrast, none of the single view models is dominant for all tasks. This demonstrates that the MMELON architecture is indeed able to leverage the “sweet spots” from each individual view to achieve better scaling across a large number of diverse downstream tasks.
For the Graph model we introduced a novel pre-training task, Betti number prediction that captures topological features. Among the single view models, the Graph Model is the strongest and is the most significant contributor to the multi-view model in most cases (Fig. 2B).
To further validate the effectiveness of MMELON we apply our pre-trained multi-view model to a large collection of GPCR assays to create over $100$ fine-tuned activity models. This serves as a virtual screening panel against which a set of metabolites and FDA approved drugs are screened. We find a large number (33) of GPCR targets that are implicated in Alzheimer’s disease. Several strong binders to these targets are identified from the metabolite and drug datasets. They are studied by molecular docking  [47] and pharmocophore identification, showing promising results. Future work will aim to further validate these predictions through wet-lab experiments.
@@ 4.3 Architectures
4.3.1 Graph model
The Graph model in our framework is based on the TokenGT architecture, which treats nodes and edges in a molecular graph as independent tokens. TokenGT provides two positional encoding schemes: orthogonal random features (ORFs) and Laplacian eigenvectors, which capture graph structure. In our implementation, we use Laplacian eigenvectors to encode positional information, as they capture the structural relationships between atoms through the graph Laplacian. The remaining layers of the model use a standard transformer architecture without graph-specific modifications. By combining explicit featurization of tokens with the self-attention mechanism of the transformer, graph model captures both local molecular features and global structural relationships without having to relearn easily computable features at token level.
4.3.2 Text model
@@ 4.4 Pre-training tasks
4.4.1 Pre-training on topological context
@@ 4.5 Late fusion strategies
4.5.1 Multi-view late fusion approach
@@
4.5.2 Alternative late fusion strategies
@@
The performance on a subset of MoleculeNet tasks for alternate schemes 1, 2, and 3 is compared against that of the multi-view model in Tables S5-S6. Overall, the chosen multi-view architecture is the top performer.
@@ 4.6 Drug-target interaction model
Ref.  [45] proposed a sequence-based framework to investigate the binding affinities of proteins and ligands whilst experimenting with various embedding types for this purpose. From this study, we adapted the model that uses convolutional neural network-based encodings obtained from the ESM-1b [43] sequences for representing the protein structure. In line with their model, we combined the CNN based protein embeddings with ligand embeddings, which were obtained from our models. The combined embeddings were then passed through a prediction head, which was again based on their proposed prediction head, expect for the difference in the size of the embeddings. We used 64-dimensional protein and ligand encodings which are then concatenated to form a 128-dimensional embedding. The authors of Ref. [45] used double this size. We experimented with combining CNN obtained protein embeddings with ligand embeddings obtained from various modalities of representation: Image, Text, Graph representations, and the combination of all three.
During training, the CNN and BMFM model weights were unfrozen gradually using the Xavier method. We used the batch size of 64 and the learning rate of 0.0005 for all models except for the Text model, for which we used the batch size of 128 and the learning rate of 0.00001. We trained for two thousand epochs, as suggested in  [45], and used the best performing model on the validation set to predict the test set. We used six-fold cross-validation using one-sixth of the data for validation, one-sixth for testing, and the rest for training. The results were then averaged across the six cross-validation folds.
@@ 4.7 Statistical Analysis of select classification models
We consider the fine-tuned output of the three single-view and multi-view models in select classification tasks and aim to perform analysis to uncover distinct characterizations of the models’ predictions. Binarization was performed on outputs by applying a threshold of 0.5 to the sigmoid mapped scores. The McNemar test [71] was applied to the predictions of model pairs to test whether one model predicts greater or less active ligands than the other model than expected by chance drawn randomly from identical distributions. The McNemar tests tend to be very sensitive to relatively “small” variations above random. $F1$ scores compare predictions, measuring how well the second model’s predicted activities predict the first model’s predicted activities. It is important to realize this measure is not symmetric.
@@ 4.8 Fine-tuning Protocols
Hyper-parameter optimization of the fine-tuning process is important to ensure quality outputs. Optimization was first explored using ray-tune [72]. Based on an exploration of our multi-view model across a subset of MoleculeNet, a set of parameters for AdamW and regularization (e.g. weight-decay) were chosen. Linear and MLP prediction heads were also explored. The learning rate was then varied for individual MoleculeNet tasks for Image, Graph, Text in the range of 1e-5 to 1e-3). In the case of the multi-view model, the presence of multiple sub-networks may call for a multiple optimizer fine-tuning strategy that employs up to four learning rates in the aforementioned range. This strategy was employed for the QM7 dataset. For CYP and GPCR models, a learning rate was chosen across the set of tasks. The highest average validation performer across 5 seeds was chosen as the model instance to report the held-out (test) set results.
In order to facilitate the plot in Fig. 2 we scale the regression metrics according to the following formula $m^{\prime}=1-(m-m_{0})/s$ where $m$ is the unscaled metric and $m_{0}$ and $s$ are an offset and scaling factor, respectively. For ESOL, FreeSolv, Lipophilicity and QM7 the parameters $(m_{0},s)$ are $(0.7,1.0)$, $(1.5,5.0)$, $(0.50,1.0)$, and $(55.0,100.0)$, respectively.
@@ 4.10 Molecular docking simulations
Top 10 drugs or metabolites of AD genetics-informed FPR1 and AD multi-omics-informed ADA2A were inspected. 3D structures of FPR1 or ADA2A were retrieved from AlphaFold2 database[73]. The preparation of 2D structures of small molecules and 3D structures of proteins refer to our previous study. Molecular docking was processed by AutoDock Vina (version 1.1.2)[74]. The top 10 binding modes were investigated, and the best binding poses were selected for further analysis. Detailed binding analysis were conducted by using PyMOL (version 3.0.3).
@@
Code Availability Our Multi-View model is available on GitHub (https://github.com/BiomedSciAI/biomed.multi-view) and Hugging Face (https://Huggingface.co/ibm/biomed.sm.mv-te-84m).
@@
Acknowledgments We acknowledge Viatcheslav Gurev, William Ogallo, Sekou Remy, Mohamed Ghalwash, Wendy Cornell, James Kozloski, Michal Rosen-Zvi and Ajay Royyuru for insightful discussions.

### v3

@@ 2.4 Case study: GPCR metabolite and drug target binding and identification
We fine-tune our multi-view model on experimental binding affinity assays for 106 GPCR targets. This dataset contains experimental pKI values curated from ChEMBL [43] and GLASS [44]. All GPCRs datasets were retrieved but the GPCRs with fewer bioactivity measurements ($<100$) were not considered. Before training, the dataset is filtered to exclude molecules with a molecular weight $>600$ Da. This cutoff is commiserate with the molecular properties of the metabolites and drug molecules that are in the screening (inference) set. Datasets with less than 50 data points were removed from the set. Each assay/GPCR target is treated as a separate regression task and data is split using the balanced scaffold protocol [25]. In Fig 3A, we plot the prediction on the held-out set against the experimental values. The multi-view fine-tuned models are overall in good agreement with the held-out set (Pearson correlation coefficient 0.78, average RMSE 0.79). To further validate our model, in Fig. 3B we plot the Pearson correlation coefficient against the RMSE for each individual target. These metrics are negatively correlated, as expected. Over $80\%$ of fine-tuned models have $\text{RMSE}<1$ and nearly $70\%$ have correlations $>0.6$ on the held-out set.
@@
Next we consider interactions of molecules in our virtual screen with specific AD-related receptor. FPR1 is suggested as a genetics-informed GPCR. High level of FPR1 is causally associated with increased risks of AD (Beta = -0.1, FDR = 0.001, Fig. S6A). Through inspecting top 10 small molecules in our screen, we found that a gut metabolite, acetyl-glutamine is predicted to interact with FPR1 (multi-view score: 6.8). By re-analyzing the abundance of gut metabolites in microbial strains in vitro [62], acetyl-glutamine was determined to have a high level in Ruminococcus gnavus (log2FoldChange: 1.19), a strain that mono-colonized in mice performed better on a spatial working memory test [63]. Additionally, over-expression of $\text{A}\beta$ protein is significantly characterized by the decreased Ruminococcus gnavus [64]. Three-dimensional (3D) binding interaction analysis revealed that acetyl-glutamine is located in an allosteric binding site that is distinct from the classical binding site, indicating a potential regulation mechanism (Fig. 3C). For top predicted drugs, glutathione (GSH), an anti-oxidant drug [65], GSH, is prioritized (multi-view score: 6.53) to interact with FPR1. GSH is a nutrition supplementation that consist of three amino acids, including glutamic acid [66]. The depletion of GSH has been found in the hippocampal regions in mild cognitive impairment (MCI) and AD patients compared to healthy control [67]. GSH forms strong hydrogen bond interactions with FPR1 in the classic binding site (Fig. 3D). We find that image ($\alpha=0.51$) and graph ($\alpha=0.39$) are the top contributors to the FPR1 multi-view model and generate attention heat maps from the image and graph sub-networks (Fig. 3C-D). Most of the functional groups of acetyl-glutamine (e.g., amide group) and GSH have high attentions. For acetyl-glutamine binding to FPR1, the image highlights the attention of amide group that forms hydrogen bond with FPR1 while the graph captures the main chain carbon that controls the direction of amide group, which is consistent with the binding mode.
@@
ADA2A is a GPCR with the strong multi-omics evidence to have a significant lower expression in AD (differential expressed in 1 human and 4 mouse transcriptomics datasets, Fig. S6B) [61]. A gut metabolite, fructose 1,6-biphosphate, is predicted to strongly interact with ADA2A (multi-view score 7.37). Fructose 1,6-biphosphate is involved in glycolysis, the level of which was down regulated in APP/PS1 mice [68]. High levels of 1,6-biphosphate have been detected in genus Mitsuokella (log2FoldChange: 13.42), which is decreased in $\text{A}\beta+$ (positive) patients[69]. Fructose 1,6-biphosphate is predicted to bind a classic site of ADA2A (Fig. 3E). Isosorbide dinitrate, a drug used to prevent and treat chest pain caused by coronary artery disease[45], is predicted as a top candidate to bind ADA2A (multi-view score: 8.21, Fig. 3F). The high attention of fructose 1,6-biphosphate and isosorbide dinitrate were captured for key chemical moieties (Fig. 3E-F). Although the image serves as the primary contributor ($\alpha=0.76$), the graph representations play a crucial supplementary role ($\alpha=0.22$) to preserve functional atoms, such as phosphorus atoms in phosphate of fructose 1,6-biphosphate and nitrogen atoms in nitro groups of isosorbide dinitrate, all of which establish key interactions with ADA2A.
@@
Figure 3: (A) Correlation plot of predicted vs. experimental pKI values in the held out set. Values of the Pearson correlation and the RMSE are given in the legend. (B) Pearson correlation coefficient plotted against RMSE for 33 Alzheimer’s disease related GPCR targets (red dots) and others (blue dots). Examples of strong binders to FPR1 (C-D) and ADA2A (E-F) are depicted in the respective panels. Binding poses generated from docking software (left column panels) are depicted alongside heat maps on attention levels as predicted by the graph (center column panels) and image (right column panels) sub-networks of our model. Darker color indicates higher attention level.
@@ 3 Discussion
A molecular foundation model that is fine-tunable on a broad, diverse set of tasks can serve as a crucial component of any discovery workflow and the choice of molecular representation is crucial. Each choice may have strengths and weaknesses, and each individual representation (view) can perform well on certain downstream tasks and not others. Here we introduce MMELON, a multi-view architecture that combines molecular representations from multiple embeddings using an interpretable late fusion approach, utilizing foundation models based on three common representations of chemical structure, Graph, Image and Text.
We fine-tune our multi-view model on a set of benchmark datasets that span diverse tasks including binding activity, drug metabolism, and ADME properties. We compare the performance of the multi-view model with the single view models whose encodings feed into it. Overall the multi-view model performs robustly across a wide range of property prediction tasks involving both classification and regression - it gives the best performance in many of the tasks, while closely matching the best single model in the remaining ones. While none of the single view models is superior for all tasks, the MMELON architecture is able to leverage the “sweet spots” from each individual view to achieve better scaling across a large number of diverse downstream tasks.
For the Graph model we introduce a novel pre-training task, Betti number prediction that captures topological features. Among the single view models, the Graph model is the strongest and is the most significant contributor to the multi-view model in most cases (Fig. 2B).
To further validate the effectiveness of MMELON we apply our pre-trained multi-view model to a large collection of GPCR assays to create over $100$ fine-tuned activity models. This serves as a virtual screening panel against which a set of metabolites and FDA approved drugs are screened. We further zoom in to a large number (33) of GPCR targets that are implicated in Alzheimer’s disease. Several strong binders to these targets are identified from the metabolite and drug datasets. They are studied by molecular docking  [58] and pharmocophore identification, showing promising results. Future work will aim to further validate these predictions through wet-lab experiments.
@@ 4.3 Architectures
Graph model
The Graph model in our framework is based on the TokenGT architecture, which treats nodes and edges in a molecular graph as independent tokens. TokenGT provides two positional encoding schemes: orthogonal random features (ORFs) and Laplacian eigenvectors, which capture graph structure. In our implementation, we use Laplacian eigenvectors to encode positional information, as they capture the structural relationships between atoms through the graph Laplacian. The remaining layers of the model use a standard transformer architecture without graph-specific modifications. By combining explicit featurization of tokens with the self-attention mechanism of the transformer, graph model captures both local molecular features and global structural relationships without having to relearn easily computable features at the token level.
Image model
Text model
@@ 4.4 Pre-training tasks
Pre-training on topological context
@@ 4.5 Late fusion strategies
Multi-view late fusion approach
@@
Alternative late fusion strategies
@@
The performance on a subset of MoleculeNet tasks for alternate schemes 1, 2, and 3 is compared against that of the multi-view model in Tables S7-S8. Overall, the chosen multi-view architecture is the top performer.
@@ 4.6 Drug-target interaction model
Ref. [56] proposed a sequence-based framework to investigate the binding affinities of proteins and ligands whilst experimenting with various embedding types for this purpose. From this study, we adapted their model for representing the protein structure that uses convolutional neural network-based encodings obtained from ESM-1b [54]. We combined these CNN-based protein embeddings with ligand embeddings obtained from our models. The combined embeddings were then passed through a prediction head, which was based on the prediction head of Ref. [56]. We used 64-dimensional protein and ligand encodings which are then concatenated to form a 128-dimensional embedding. The authors of Ref. [56] used double this size. We experimented with combining the protein embeddings with ligand embeddings obtained from various modalities of representation developed here: Image, Text, Graph, and multi-view.
During training, the protein, single- and multi- view model weights were unfrozen gradually using the Xavier method. We used a batch size of 64 and a learning rate of 0.0005 for all models except for the Text model, for which we used a batch size of 128 and a learning rate of 0.00001. We trained for 2000 epochs, as suggested in Ref [56], and used the best performing model on the validation set to predict the test set. We used six-fold cross-validation using one-sixth of the data for validation, one-sixth for testing, and the rest for training. The results were then averaged across the six cross-validation folds.
@@ 4.7 Statistical analysis of select classification models
We consider the fine-tuned output of the three single-view and multi-view models in single-task classification and regression tasks and aim to perform comparative analyses to uncover distinct characterizations of the models’ predictions. For classification tasks, binarization was performed on outputs by applying a threshold of 0.5 to the sigmoid mapped scores. $F1$ scores are employed to evaluate the agreement between models. $F1$ scores compare binarized predictions typical of classifications, measuring how well the second model’s predicted activities predict the first model’s predicted activities. Note that this measure is not symmetric. For regression tasks, the model outputs for continuous prediction results are compared using the Pearson correlation coefficient, with a $\chi^{2}$ p-value estimator.
@@ 4.8 Fine-tuning protocols
The choice of data splitting into training, validation and test sets greatly impacts model performance. For example, The size-ordered scaffold split clusters molecules according to their Murko scaffold [81] and assigns the clusters to splits in size-order. This is to be compared with other variations of scaffold split, which cluster the dataset in the same manner but assign clusters to train, validation, and test partitions randomly or ‘balanced’ (see, e.g. Ref.  [25]). The balanced scaffold split assigns clusters large relative to the test set size to the training and other clusters randomly. Scaffold splits are generally understood to be more appropriate and harder than random splits, though the latter are still often utilized in benchmarks [34]. Following Ref.  [25] size-ordered scaffold and were used for MoleculeNet and balanced scaffold splits for CYP and GPCR tasks with an 0.8/0.1/0.1 train/validation/test splitting. For the ComputationalADME datasets it is noted that scaffold clustering yields a large number of singletons [49] so a random split (0.7/0.1/0.2 train/validation/test) was employed.
Hyper-parameter optimization of the fine-tuning process is important to ensure quality outputs. Optimization was first explored using Ray-Tune [82]. Based on an exploration of our multi-view model across a subset of MoleculeNet, a set of parameters for AdamW and regularization (e.g. weight-decay) were chosen. Linear and MLP prediction heads were also explored. The learning rate was then varied for individual MoleculeNet tasks for Image, Graph, Text in the range of 1e-5 to 1e-3). In the case of the multi-view model, the presence of multiple sub-networks may call for a multiple optimizer fine-tuning strategy that employs up to four learning rates in the aforementioned range. For CYP, ComputationalADME and GPCR models, a learning rate was chosen across each set of tasks. The highest average validation performer across 5 seeds was chosen as the model instance to report the held-out (test) set results.
In order to facilitate the plot in Fig. 2 we scale the regression metrics according to the following formula $m^{\prime}=1-(m-m_{0})/s$ where $m$ is the unscaled metric and $m_{0}$ and $s$ are an offset and scaling factor, respectively. For ESOL, FreeSolv, Lipophilicity and QM7 the parameters $(m_{0},s)$ are $(0.7,1.0)$, $(1.5,5.0)$, $(0.50,1.0)$, and $(55.0,100.0)$, respectively. For the six ComputationalADME datasets, parameters $(0.30,1.0)$ are used.
@@ 4.10 Molecular docking simulations
3D structure of CYP2C9 complexed with a drug Lorsatan was retrieved from Protein Data Bank (PDB ID: 8VX0)[53]. The top bioactive molecules (CHEBL1466628, CHEMBL1353420, and CHEMBL1742011) were prioritized from our classification model. Top 10 drugs or metabolites of AD genetics-informed FPR1 and AD multi-omics-informed ADA2A were inspected. 3D structures of FPR1 or ADA2A were retrieved from AlphaFold2 database[83]. The preparation of 2D structures of small molecules and 3D structures of proteins refer to our previous study[58]. Molecular docking was processed by AutoDock Vina (version 1.1.2)[84]. Heme was kept for CYP2C9 molecular docking. The top 10 binding modes were investigated, and the best binding poses were selected for further analysis. Detailed binding analysis were conducted by using PyMOL (version 3.0.3).
@@
Code Availability Our Multi-View model is available on GitHub (https://github.com/BiomedSciAI/biomed-multi-view) and Hugging Face (https://Huggingface.co/ibm/biomed.sm.mv-te-84m).
@@
Acknowledgments We acknowledge Viatcheslav Gurev, William Ogallo, Sekou Remy, Mohamed Ghalwash, Leili Zhang, Wendy Cornell, James Kozloski, Michal Rosen-Zvi and Ajay Royyuru for insightful discussions.